# Moonworks Lunara Aesthetic II: An Image Variation Dataset

### v1

@@ Abstract
Nada SolimanCarnegie Mellon University, nsoliman@alumni.cmu.edu, M M Sayeef Abdullah and Sabit Hassan
@@ 1.  Introduction
Recent frontier text-to-image systems, including OpenAI’s GPT-4o (OpenAI (2025b, a)) and Google’s Gemini image models (Google (2025)), exhibit substantial advances in instruction following, compositional reasoning, and visual consistency. The outputs of such systems, however, are typically governed by restrictive usage terms that prohibit their use for training or developing competing models. While recent open-source efforts such as Z-image (Team (2025b)) and Qwen-Image/Edit (Wu et al. (2025)) represent important progress, the field continues to lack high-quality, high-aesthetic training data at sufficient signal density. As such, we publicly release Lunara Aesthetic II, a highly curated dataset comprising 2.8K anchor-linked variation sets generated from original artwork, photographs, and the Moonworks Lunara model. The dataset is explicitly designed to operationalize identity-preserving contextual variation, enabling image generation and editing systems to be trained and evaluated on genuine contextual generalization rather than surface-level visual memorization.
@@
In contrast to large-scale image–text datasets such as LAION-5B Schuhmann et al. (2022a), which prioritize scale and coverage, and web-harvested caption datasets such as Conceptual Captions and YFCC100M Sharma et al. (2018); Thomee et al. (2016), which lack explicit supervision for identity or contextual consistency, Lunara Aesthetic II is explicitly constructed around anchor-linked, identity-preserving variation. Instruction-based image editing benchmarks, including InstructPix2Pix and MagicBrush Brooks et al. (2023); Zhang et al. (2023), introduce text-guided edits and largely consist of isolated, single-step transformations. Lunara Aesthetic II is designed to unify high aesthetic quality with structured, identity-preserving contextual variation, enabling principled training and evaluation of controlled generalization in image generation and editing models.
@@ 2.3 Conditional Co-occurrence of contextual Variations
(f) Change of weather and viewpoint camera
Figure 4: Contextual variations in original photographs. Each pair shows (a) and (b).
@@ 3.  Dataset Construction
All original images were obtained with explicit contributor consent. We first distilled initial image prompts using QWEN3-VL Team (2025a), then generated 3,324 candidate variations with Moonworks Lunara, a diffusion mixture model that is trained using Moonworks CAT (Composite Active Transfer)111further details of Lunara and CAT will be provided in upcoming release, a pretraining-time approach inspired by active learning (Hassan et al. (2025a, b, 2024); Hassan and Alikhani (2023); Hassan et al. (2018)).
@@ 4.  Dataset Evaluations
(a) Change of scene-composition
(b) Change of viewpoint camera + color-tone
(c) Change of illumination time and mood
(d) Change of scene-composition and viewpoint camera
Figure 5: contextual variations in original art. Each pair shows (a) and (b).
@@ 4.2 Aesthetic Preference Analysis
We evaluate image aesthetics using the LAION Aesthetics v2 predictor, a CLIP-based model trained to approximate aggregate human judgments of visual appeal. We compare Lunara-AIV against several widely used vision–language datasets, including Lunara Art Aesthetic I Wang et al. (2026), Conceptual Captions (CC3M) Sharma et al. (2018), a random subset of LAION-2B-Aesthetic Schuhmann et al. (2022b), and the Wikipedia-based Image–Text dataset (WIT) Srinivasan et al. (2021). Table 3 reports full distributional statistics of predicted aesthetic scores.
@@ 4.4 Evaluation results
Human evaluation results indicate that the generated image variations maintain a high degree of identity stability, with a mean Likert score of 4.68 across all evaluated samples. In terms of target attribute realization, performance is consistently strong across contextual dimensions, achieving 93.3% accuracy for illumination changes and 90.5% for weather-related variations. Viewpoint changes are correctly realized in 88.2% of cases, while object composition modifications reach 84.0% accuracy. More subtle appearance-related deltas show slightly lower but still robust performance, with 85.8% accuracy for color tone and 81.5% for mood or atmospheric changes. Overall, the results suggest that the model reliably implements a wide range of contextual transformations while preserving the core identity of the scene.
@@ 5.  Discussion
Lunara Aesthetic II is intentionally released separately from the Lunara Aesthetics Dataset Wang et al. (2026). While the Aesthetics dataset targets stylistic conditioning and aesthetic control through broad prompt and artistic diversity, Lunara Aesthetic II focuses on contextual variation, controlled semantic transformation, and identity preservation under change. This separation clarifies intended use and evaluation: improvements on Lunara primarily reflect advances in style modeling, whereas improvements on Lunara Aesthetic II more directly measure contextual consistency and transformation behavior while maintaining high visual quality.
@@ 6.  Conclusion and Future Release
From an ethical perspective, all images are derived from original, ethically sourced content and curated to avoid personally identifiable information (PII).. We encourage responsible use aligned with the dataset’s purpose: controlled evaluation of contextual transformations and identity-preserving edits.

### v3

@@ Abstract
Nada SolimanCarnegie Mellon University, nsoliman@alumni.cmu.edu, Sayeef Abdullah and Sabit Hassan
@@ 1.  Introduction
Recent frontier text-to-image systems, including OpenAI’s GPT-4o (OpenAI (2025a, b)) and Google’s Gemini image models (Google (2025)), exhibit substantial advances in instruction following, compositional reasoning, and visual consistency. The outputs of such systems, however, are typically governed by restrictive usage terms that prohibit their use for training or developing competing models. While recent open-source efforts such as Z-image (Team (2025b)) and Qwen-Image/Edit (Wu et al. (2025)) represent important progress, the field continues to lack high-quality, high-aesthetic training data at sufficient signal density. As such, we publicly release Lunara Aesthetic II, a highly curated dataset comprising 2.8K anchor-linked variation sets generated from original artwork, photographs, and the Moonworks Lunara model. The dataset is explicitly designed to operationalize identity-preserving contextual variation, enabling image generation and editing systems to be trained and evaluated on genuine contextual generalization rather than surface-level visual memorization.
@@
In contrast to large-scale image–text datasets such as LAION-5B (Schuhmann et al. (2022a)), which prioritize scale and coverage, and web-harvested caption datasets such as Conceptual Captions and YFCC100M (Sharma et al. (2018); Thomee et al. (2016)), which lack explicit supervision for identity or contextual consistency, Lunara Aesthetic II is explicitly constructed around anchor-linked, identity-preserving variation. Instruction-based image editing benchmarks, including InstructPix2Pix and MagicBrush (Brooks et al. (2023); Zhang et al. (2023)), introduce text-guided edits and largely consist of isolated, single-step transformations. Lunara Aesthetic II is designed to unify high aesthetic quality with structured, identity-preserving contextual variation, enabling principled training and evaluation of controlled generalization in image generation and editing models.
@@ 2.3 Conditional Co-occurrence of contextual Variations
Hogsmede houses with snow.
Hogsmede in Christmas.
@@
Sunset on lake with lilies.
Lake at night with lilies.
@@
Serene lake with boats..
Serene lake in autumn..
@@
Flower vase on table..
Flower vase near window..
@@
Christmas trees in park..
Christmas trees in snow..
@@
Looking over a lake..
From further away, in fall..
(f) Change of weather
Figure 4: Contextual variations in original photographs. Each pair shows original (left) and modified image (right).
two figures and wolves..
a figure and two wolves..
(a) Change of scene-composition
bicycle in front of..
bicycle around corner..
(b) Change of viewpoint camera + color-tone
next to a goldfish bowl..
warm light and mood..
(c) Change of illumination time and mood
hiker on snowy mountain..
walking down to a river..
(d) Change of scene-composition
Figure 5: Contextual variations in original art. Each pair shows original (left) and modified images (right).
@@ 3.  Dataset Construction
All original images were obtained with explicit contributor consent. We first distilled initial image prompts using QWEN3-VL (Team (2025a)), then generated 3,324 candidate variations with Moonworks Lunara, a diffusion mixture model that is trained using Moonworks CAT (Composite Active Transfer)111further details of Lunara and CAT will be provided in upcoming release, a pretraining-time approach inspired by active learning (Hassan et al. (2025a, b, 2024); Hassan and Alikhani (2023); Hassan et al. (2018)).
@@ 4.2 Aesthetic Preference Analysis
We evaluate image aesthetics using the LAION Aesthetics v2 predictor, a CLIP-based model trained to approximate aggregate human judgments of visual appeal. We compare Lunara-AIV against several widely used vision–language datasets, including Lunara Art Aesthetic I (Wang et al. (2026)), Conceptual Captions (CC3M) (Sharma et al. (2018)), a random subset of LAION-2B-Aesthetic (Schuhmann et al. (2022b)), and the Wikipedia-based Image–Text dataset (WIT) (Srinivasan et al. (2021)). Table 3 reports full distributional statistics of predicted aesthetic scores.
@@ 4.3.1 Human Evaluation results
Results indicate that the generated image variations maintain a high degree of identity stability, with a mean Likert score of 4.68 across all evaluated samples. In terms of target attribute realization, performance is consistently strong across contextual dimensions, achieving 93.3% accuracy for illumination changes and 90.5% for weather-related variations. Viewpoint changes are correctly realized in 88.2% of cases, while object composition modifications reach 84.0% accuracy. More subtle appearance-related deltas show slightly lower but still robust performance, with 85.8% accuracy for color tone and 81.5% for mood or atmospheric changes. Overall, the results suggest that the model reliably implements a wide range of contextual transformations while preserving the core identity of the scene.
@@ 5.  Discussion
Lunara Aesthetic II is intentionally released separately from the Lunara Aesthetics Dataset (Wang et al. (2026)). While the Aesthetics dataset targets stylistic conditioning and aesthetic control through broad prompt and artistic diversity, Lunara Aesthetic II focuses on contextual variation, controlled semantic transformation, and identity preservation under change. This separation clarifies intended use and evaluation: improvements on Lunara primarily reflect advances in style modeling, whereas improvements on Lunara Aesthetic II more directly measure contextual consistency and transformation behavior while maintaining high visual quality.
@@ 6.  Conclusion and Future Release
From an ethical perspective, all images are derived from original, ethically sourced content and curated to avoid personally identifiable information (PII). We encourage responsible use aligned with the dataset’s purpose: controlled evaluation of contextual transformations and identity-preserving edits.